Solve the system 2x + 3y = 18
4x - 5y = 14

Answer by mananth(16946)

2x + 3y = 18.............1
4x - 5y = 14 .............2
multiply equation1 by 2 and subtract eq2
4x+6y-4x+5y=36-14
11y=22
y=2
..
plug the value of y in equation 1
2x+3y=18
2x+6=18
2x=18-6
2x=12
x=6